OpenStudy (anonymous):

There are 3 inlet taps and 1 outlet tap with diameters 8, 6, 2 (cms – inlet taps) and 4 cm respectively. The rate of flow of oil is directly proportional to the square of the radius of the tap. It takes 9 minutes for the smallest tap to fill an empty tank. What is the time taken to fill an empty tank when all the taps (inlet + outlet) are kept open?

OpenStudy (anonymous):

shoot i just answered this and lost it somehow!! i've no time to type it now = the answer i got was 9/22 minutes

OpenStudy (anonymous):

scroll down for related answers in the link I provided. You should get an idea as to how to approach this problem

OpenStudy (anonymous):

could you help how to calcuate the rate of flow

OpenStudy (anonymous):

rate of flow is proportional to radius squared so flow rate of tap A = k*8^2 = 64k similarly the flow rates of the other three are 36k, 4k and 16k

OpenStudy (anonymous):

Hi, but shouldn't the radius be 4, since diameter is 8

OpenStudy (anonymous):

oh you are right. well spotted. so it is 16k, 9k, k and 4k

OpenStudy (anonymous):

getting the answer as 18/22, but it seems to be 7/22

OpenStudy (anonymous):

total flow = 26-4= 22

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

total flow rate is 22k per minute. so what is the next step? what is the value of k?

OpenStudy (anonymous):

total capacity of the tank is 36 units (2^2*4.5)

OpenStudy (anonymous):

total flow that comes is 16+9+1-4==22

OpenStudy (anonymous):

ok got it, total 9 units, 9/22

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

np. glad you could work it out on your own.